# Security Weaknesses of Copilot Generated Code in GitHub

### v2

@@ 38 Types Total: 544
2 CWE-78 (Improper Neutralization of Special Elements used in an OS Command) CWE-95 (Improper Neutralization of Directives in Dynamically Evaluated Code)
3 CWE-772 (Missing Release of Resource after Effective Lifetime) CWE-563 (Assignment to Variable without Use)
4 CWE-89 (Improper Neutralization of Special Elements used in an SQL Command) CWE-20 (Improper Input Validation)
5 CWE-259 (Use of Hard-coded Password) CWE-185 (Incorrect Regular Expression)
4.3. RQ3: How many security weaknesses belong to the CWE Top-25?
Approach. The code snippets in our collected dataset were mainly generated in 2022 and 2023. To compare whether the security issues in Copilot-generated code are widespread in this period, we chosed MITRE 2023 CWE Top-25 list (The MITRE Corporation, 2023a) as our baseline. In addition, MITRE maintains another list of “Stubborn Weaknesses in the CWE Top 25” (The MITRE Corporation, 2023d), which includes those CWEs that have consistently appeared in the Top 25 most dangerous software weaknesses of the past five years (2019-2023). We compared the CWEs obtained in RQ2 with the 2023 CWE Top 25 and “Stubborn Weaknesses” in the CWE Top 25.
Results. The distribution of CWEs found compared to the MITER list is shown in Table 7. The results show that the CWE weaknesses present in the code generated by Copilot belong to eight CWEs. We highlight those CWEs that are “Stubborn Weaknesses” (SW) in bold. This means these are present issues and are currently among the most common and serious security weaknesses in practice. It is worth noting that the 218 security issues present in the code snippet correspond to these eight CWEs, while another 30 CWEs cover the remaining 326 security issues. This indicates that the CWE Top-25 weaknesses are also prevalent in the code generated by Copilot. At the same time, we can see that CWE-78: OS Command Injection occurs more frequently in Copilot-generated code. CWE-94: Improper Control of Generation of Code is in the list of the CWE Top-25 (The MITRE Corporation, 2023a), and meanwhile it is also one of the weaknesses with a high occurrence frequency in our results. Some CWEs with a higher ranking in the Top-25 list do not frequently appear in Copilot-generated code, such as CWE-79: Cross-site Scripting and CWE-89: SQL Injection.
Table 7. The CWEs that belong to the 2023 CWE Top-25 list
CWE-ID Description Is SW # Related Snippets Frequency
CWE-94 Improper Control of Generation of Code No 34 115
CWE-78 Improper Neutralization of Special Elements used in an OS Command Yes 18 81
CWE-798 Use of Hard-coded Credentials Yes 3 6
CWE-502 Deserialization of Untrusted Data Yes 2 6
CWE-20 Improper Input Validation Yes 3 6
CWE-79 Improper Neutralization of Input During Web Page Generation Yes 2 2
CWE-276 Incorrect Default Permissions No 1 1
@@ 4.3. RQ3: How many security weaknesses belong to the CWE Top-25?
CWE-22 Improper Limitation of a Pathname to a Restricted Directory Yes 1 1
Total 218
@@ 5. Discussion
In this section, we explain the study results in Section 5.1 and then discuss their implications in Section 5.2.
@@ 5.1. Interpretation of Results
RQ1: How secure is the code generated by Copilot?
@@
Among the 452 code snippets generated by Copilot, we found that 29.6% of these code snippets contain security weaknesses. Although in our results, the proportion of security issues in Python code snippets is higher than in JavaScript, there is no significant difference. Using Copilot to write code in Python or JavaScript can generally lead to security issues. This could be attributed to features that made their code more flexible, such as dynamic typing and interpretation. From a Copilot point of view, the generated code does not need to reason about the whole program because it is dynamically typed, consequently, it does not need to read the flow of the whole program to generate working suggestions. In summary, developers should pay special attention to the security of Copilot-generated code, taking appropriate measures to validate input data and manage resources effectively to minimize security risks. The results of RQ1 suggest that in practical production, Copilot can help developers write code faster and increase productivity (GitHub, 2024; Perry et al., 2022). Still, additional security assessments and fixes are also required to ensure that the generated code does not introduce potential security risks.
RQ2: What security weaknesses are present in the code snippets generated by GitHub Copilot?
@@
After conducting a security evaluation of 452 code snippets generated by Copilot, a total of 544 security weaknesses were identified, involving 38 CWEs, which is around 10% of the CWEs (439 CWEs) (The MITRE Corporation, 2023c). This may be because Copilot generates code in different programming languages and application scenarios. In addition, since the Copilot base model (Codex) is trained on publicly available data that potentially contain various types of security weaknesses, this can lead to the presence of multiple CWEs in the generated code by Copilot. This set of 38 CWEs covers many security issues, and Table 8 shows the types of security issues to which these CWEs are relevant.
Table 8. CWEs and security issue types in Copilot-generated code
Type of Security Issue Relevant CWEs
Data Neutralization Issues CWE-94, CWE-78, CWE-117, CWE-89
Resource Management Errors CWE-772, CWE-502
Error Conditions, Return Values CWE-396, CWE-617
Bad Coding Practices CWE-628, CWe-563
Credentials Management Errors CWE-798
Information Management Errors CWE-312, CWE-215
Authentication Errors CWE-295
Concurrency Issues CWE-367
File Handling Issues CWE-22
The diversity of security weaknesses indicates that developers using Copilot face various security risks. These risks are diverse, covering different development environments and application scenarios. In addition, we find that CWEs in Python are mainly related to data processing and system calls. In contrast, CWEs in JavaScript are often associated with dynamic code generation problems and security issues in web development. This may be because the two languages are designed differently. JavaScript is more commonly used for web development, while Python is more widely used in areas such as data processing and scientific computing. Overall, while Python and JavaScript differ in some common types of security weaknesses, they require developers to be aware of and take timely and targeted security measures to mitigate these risks. For example, developers should perform adequate validation of user inputs. It is also necessary to restrict the program’s permissions so that they only access essential resources.
The results of RQ2 reveal the security weaknesses that developers may encounter in an actual production environment and their frequency of occurrence, which can help developers to be aware of the security aspects of the code generated by Copilot and to take appropriate measures to address the security weaknesses in an informed manner.
RQ3: How many security weaknesses belong to the CWE Top-25?
@@
As shown in Table 7, eight of the CWEs in the Copilot-generated code can be found in the 2023 CWE Top-25 list, covering more than 218 security issues (40.1% of 544 identified CWEs) in our dataset. This indicates that the commonly acknowledged CWE Top-25 weaknesses in software development, which are considered the most prevalent and dangerous, are also prevalent in the code generated by Copilot. Therefore, developers using Copilot must pay close attention to these weaknesses and take appropriate measures to prevent them before they are integrated into their code. Meanwhile, we found that some higher-ranked CWEs in the CWE Top-25 list, such as CWE-89: SQL Injection, did not frequently appear in our scan results. This could be because the generated code does not establish connections with external resources (e.g., databases). Besides, we also observed that some vulnerabilities from the CWE Top-25 list were not detected in our scans, indicating that Copilot may sanitize and prevent specific weaknesses from being suggested to developers. GitHub is gradually improving the security of Copilot and its underlying model (Codex) (GitHub, 2023). However, it has also been found that Copilot’s security layer can potentially handle some CWEs better than others, leaving applications vulnerable to some critical vulnerabilities (Majdinasab et al., 2024).
@@
We also identified 30 security weaknesses in the code that do not belong to the CWE Top-25 list. Although these less common security weaknesses may not be as widespread as CWE Top-25, attackers can still exploit them. For example, we only detected one instance of CWE-732: Incorrect Permission Assignment for Critical Resource in our dataset. This security weakness is not commonly found in code and only occurs when specific users have certain permissions. However, it can lead to significant security risks when it does occur. Developers should also be aware of these less common security weaknesses to fully protect their code from attacks.
@@ 5.2. Implications
Importance of continuous security analysis of Copilot-generated code. We conjecture practitioners using Copilot will likely encounter security weaknesses, regardless of the programming language used. Our study results reflect the inevitability of security weaknesses in Copilot-generated code, and there are a variety of security weaknesses in Copilot-generated code. Practitioners must be aware of the diverse security scenarios in production and adopt multiple security prevention measures to address security risks before accepting vulnerable code suggestions. Maintaining a rigorous process of security checks concurrently with code generation can identify potential vulnerability risks and rectify the security issues in time. Developers should follow the best practices for using code generation tools and always check the code suggestions generated by Copilot (or any code generation tools). For example, developers can establish a gated check-in build process for checking and preventing security issues when committing code generated by Copilot. Initially, we can turn to automated tools to continuously scan the Copilot-generated code for known security issues, such as CWEs. Recognizing that these automated tools may not always detect all security issues, especially newly emerging ones or those that are context-specific, it is necessary to conduct a subsequent manual assessment, including manual security code review for Copilot-generated code. By embracing this combined strategy for continuous security analysis, we can ensure a robust security shield for the code committing process with Copilot-generated code.
Prevention of security issues in Copilot-generated code. According to our study results, we provide the following suggestions on how to prevent potential security issues in Copilot-generated code: (1) Targeted security countermeasures: Based on the frequency of related CWEs in Copilot-generated code, practitioners can proactively prevent and address security issues in a targeted manner. When using Copilot to generate Python or JavaScript code, developers should focus on dedicated security weaknesses (CWEs) in Copilot-generated code of different programming languages (see Table 6). Furthermore, a recent study shows that security weaknesses appear in certain code suggestions, but not all (Majdinasab et al., 2024); consequently, developers should carefully select potentially more secure suggestions, with the assistance of tools, that do not expose the code to vulnerabilities. (2) Standardized security assessment: Common security weaknesses in software development are also prevalent in the code generated by Copilot. As a good practice, developers can use the CWE Top-25 list as a guide to understand which security weaknesses are most common and dangerous in the generated code and follow the mitigation measures for related CWEs provided by the MITRE (The MITRE Corporation, 2023b). Additionally, the CWE Top-25 provides a standardized approach for security assessment, and developers can also use it as a guide to perform security audits of the code generated by Copilot. (3) Enhancing prompt engineering to generate security code: The instruction tuning schemes in code generation not only impacts the utility of the code but also its security. By combining the security fine-tuning with standard instruction tuning, joint optimization of security and utility can be promoted (He et al., 2024). We should incorporate security considerations from the initial stages of code generation. For specific security scenarios that are prone to CWEs, we can improve the security of code by enhancing prompt engineering, e.g., with the prompt patterns proposed in (White et al., 2023).

### v3

@@ 43 Types Total: 628
2 CWE-78 (Improper Neutralization of Special Elements used in an OS Command) CWE-79 (Improper Neutralization of Directives in Dynamically Evaluated Code)
@@
3 CWE-427(Uncontrolled Search Path Element) CWE-22 (Improper Limitation of a Pathname to a Restricted Directory (’Path Traversal’))
4 CWE-457 (Use of Uninitialized Variable) CWE-685 (Function Call with Incorrect Number of Arguments)
5 CWE-772 (Missing Release of Resource after Effective Lifetime) CWE-770 (Allocation of Resources Without Limits or Throttling)
4.3. RQ3: Can Copilot Chat fix the security weaknesses of code generated by Copilot?
Approach. To answer RQ3, we conducted experiments using Copilot Chat on the randomly selected code snippets containing security weaknesses and asked Copilot Chat to provide fix suggestions based on the three prompts (see Section 3.4). We then fixed the code according to the suggestion provided by Copilot Chat. Specifically, we used Copilot Chat to repair snippets with security weaknesses. The repaired code snippets are then checked for security weaknesses using the same static analysis tools. After getting the analysis results, The first author used the same filtering steps described in Section 3.3.2 to filter the results. We first removed the results that were repeatedly reported by two of the tools. Then, we removed the analysis results that were not security weaknesses. Among the three types of warnings from CodeQL analysis, we only counted code snippets that had warnings and errors, and ignored recommendations on code quality. Since the code snippets we selected all come from the Repo label, it is easier to filter the results related to security weaknesses, as we did not need to locate the specific line numbers of the code generated by Copilot in each file. Based on the results of the two successive security analysis, we determine whether the security weakness was successfully removed and the code is fixed.
Results. Table LABEL:table:fix shows the proportion of CWEs that were fixed under the three prompts. We can see that using the slash command fix can fix 19.3% of security weaknesses, while our designed basic and enhanced prompts can fix 31.8% and 55.5% of security weaknesses, respectively. Copilot Chat demonstrates promising capability in fixing security weaknesses regardless of the prompt used. At the same time, providing enhanced information in the prompts helps improve the proportion of fixes. Figure LABEL:fig:cwe-repair shows the number of security weaknesses fixed in code snippets of different languages when using various prompts. With basic prompts, security weaknesses were fixed in 29.4% (48 of 163) of Python code and 34.8% (46 of 132) in JavaScript. When using the enhanced prompt, the percentage of security weaknesses fixed in both languages rose to 58.3% (95 of 163) and 51.5% (68 of 132), respectively. We can see that Copilot Chat’s ability to fix security weaknesses in Python and JavaScript snippets does not differ significantly.
Prompt Total CWE Before Fix Total CWE After Fix Fix Rate (%)
/fix 295 238 19.3%
basic 295 201 31.8%
enhanced 295 132 55.5%
Furthermore, Figure LABEL:fig:cwe-fix shows a heatmap with the number of different CWEs that were fixed when using different prompts. We can visually observe that Copilot Chat’s effectiveness in fixing different CWEs varies. For certain CWEs, Copilot Chat can perform well in repairing those snippets and provide a complete fix (i.e., remove the CWE from the code) such as CWE-259: Use of Hard-coded Password and CWE-685: Function Call with Incorrect Number of Arguments, which can be 100% fixed with the enhanced prompt. However, for some security weaknesses, the repair performance of Copilot Chat is considerably poor as it fails to aviod certain CWEs. For example, for CWE-94: Code Injection, Copilot Chat’s fixes were successful in less than 20% of cases, and Copilot Chat was unable to fix CWE-78: OS Command Injection. In addition, we also observe that as the information provided in the prompts increases, the percentages of certain CWEs being fixed improve, such as CWE-79: Cross-site Scripting and CWE-330: Use of Insufficiently Random Values Weakness. The percentage of successful fixes for CWE-79: Cross-site Scripting increases from 0% to 70%, while the percentage of successful fixes for CWE-330: Use of Insufficiently Random Values Weakness increases from less than 10% to 98%.
@@ 5.1. Interpretation of Results
RQ1: How secure is the code generated by Copilot in GitHub projects?
@@
Around quarter of the code snippets in our dataset contained security weaknesses. We found that the code with security weaknesses in Python primarily involves Utility Tools and Game applications, whereas, in JavaScript, Web Application code constitutes the majority of security weaknesses followed by Utility Tools. This disparity may arise from the two languages’ differing purposes and design philosophies (Reichenbach, 2022; Halfnine, 2024). JavaScript is mainly used for front-end development and Web applications (PubNub, 2024), leading to its security weaknesses being concentrated in scenarios related to the web and user interaction. In contrast, Python is more broadly applied in fields such as data processing and scientific computing (Nagpal and Gabrani, 2019; Simplilearn, 2024), resulting in security weaknesses occurring across a wider range of application domains. Although our results show that the proportion of security weaknesses in Python code snippets is higher than in JavaScript, there is no significant difference between the two languages. Using AI code generation tools to write code in Python or JavaScript can generally lead to security weaknesses. This could be attributed to features that made their code more flexible, such as dynamic typing and interpretation. From a tool point of view, the generated code does not need to reason about the whole program because it is dynamically typed, and consequently, it does not need to read the flow of the whole program to generate working suggestions. In summary, developers should pay careful attention to the potential security weaknesses present in automatically generated code, taking appropriate measures to validate input data and manage resources effectively to minimize security risks. RQ1 results suggest that using Copilot is prone to security issues in production and additional security assessments are required to ensure the generated code does not introduce potential security risks.
@@
RQ2: What security weaknesses are present in the code generated by Copilot across different application domains? After conducting a security analysis of 733 code snippets generated by Copilot and other tools, a total of 628 security weaknesses were identified, involving 43 CWEs, which is around 10% of the CWEs (439 CWEs) (The MITRE Corporation, 2023). This may be because Copilot and the other two tools generate code in different programming languages and application scenarios. In addition, since the base models of AI code generation tools (such as Codex) are trained on publicly available data that potentially contain various types of security weaknesses, this can lead to the presence of multiple CWEs in the generated code. This set of 43 CWEs covers many security weaknesses.
@@
The diversity of security weaknesses indicates that developers using Copilot face various security risks. These risks are diverse, covering different development environments and scenarios. At the same time, eight of the CWEs identified in our dataset can be found in the 2023 CWE Top-25 list, covering more than 200 security weaknesses (37.1% of 628 identified). This indicates that the commonly acknowledged Top-25 CWEs, which are considered the most prevalent and dangerous weaknesses, are also prevalent in the AI-generated code. Therefore, developers using Copilot must pay close attention to these weaknesses and take appropriate measures to prevent them before they are integrated into their code base. Meanwhile, we observed that some vulnerabilities from the CWE Top-25 list were not detected in our analysis, indicating that Copilot may sanitize and prevent specific weaknesses from being suggested to developers.
@@
We also identified 30 security weaknesses in the generated code that do not belong to the CWE Top-25 list. Although these are less common security weaknesses and may not be as widespread as CWE Top-25, attackers can still exploit them. For example, we only detected one instance of CWE-732: Incorrect Permission Assignment for Critical Resource in our dataset. This security weakness is not commonly found in code and only occurs when specific users have certain permissions. However, it can lead to significant security risks when it does occur.
Besides, we observed that the types of CWEs in generated code are closely related to the application domains. Code snippets used in Game development are prone to issues such as CWE-330: Use of Insufficiently Random Values Weakness and CWE-284: Improper Access Control. This may be due to the complex logic and player inputs typically involved in game applications, which can lead to security problems related to memory management and input validation. In the Utility Tool category, the proportion of security weaknesses is relatively high, which may be attributed to the diversity of user inputs. Furthermore, we found that Python code snippets often exhibit security weaknesses such as CWE-330: Use of Insufficiently Random Values Weakness, CWE-427: Uncontrolled Search Path Element, and CWE-78: OS Command Injection, whereas JavaScript code snippets primarily encounter issues like CWE-22: Path Traversal and CWE-94: Code Injection. This may be related to the differences in the application characteristics of the two languages: JavaScript, particularly through environments like Node.js, is use for file operations and dynamic code execution (Chhetri, 2016), while Python tends to focus more on system-level operations and automation scripts. In addition, the measures we need to take are different: Utility Tools in JavaScript need to emphasize path validation and the security of dynamic execution to prevent attacks such as path traversal and code injection, while those in Python should focus on ensuring safe library loading paths and validating user input for command execution to prevent system-level vulnerabilities caused by malicious input. In Web development, Python is commonly used for frequent interactions between the back-end and the database (Simplilearn, 2024), which makes it prone to CWE-89: SQL Injection. CWE-89: SQL Injection typically occurs when user input is directly embedded into SQL query statements without parameterized queries or input validation. On the other hand, JavaScript, as a front-end language that interacts directly with users, often faces security weaknesses such as CWE-79: Cross-site Scripting, CWE-94: Code Injection, and CWE-22: Path Traversal. This may be attributed to the frequent interactions between users and data in Web development. Additionally, the complexity of dependencies and the handling of sensitive data further increase the likelihood of security weaknesses.
Regarding the Network Communication category, the proportion of security weaknesses is relatively lower than other categories. This may be attributed to the fact that code in the Network Communication domain typically demands a high level of security, benefiting from mature protocols, stringent industry standards. Additionally, the protocols involved in network communication inherently incorporate certain security mechanisms to safeguard the data transmission process, further mitigating common attack risks.
Overall, while Python and JavaScript differ in some common types of security weaknesses, they require developers to be aware of and take timely and targeted security measures to mitigate these risks. RQ2 findings note the security weaknesses that developers may encounter in an actual production environment and their occurrence frequency. These findings can help developers be aware of the security aspects of the code generated by AI code generation tools and take appropriate measures to address the security weaknesses in an informed manner.
RQ3: Can Copilot Chat help fix the security weaknesses found in the code generated by Copilot? As shown in Table LABEL:table:fix, we can see that regardless of the prompts used, Copilot Chat can fix a certain proportion of security weaknesses in the code we analyzed. Even the slash command /fix can resolve nearly 20% of security problems, indicating that Copilot Chat has some capability to fix certain weaknesses. Copilot Chat is an enhanced feature that extends the capabilities of the original Copilot (GitHub, 2023). In fact, the security of Copilot and its underlying model (Codex) is continually improving (GitHub, 2023). However, it has also been found that Copilot’s security layer can potentially handle some CWEs better than others, leaving applications vulnerable to some critical vulnerabilities (Majdinasab et al., 2024). It is believed that similar programming-based chatbots, such as Amazon Q (which works with CodeWhisperer), also offer similar repair capabilities. Furthermore, we can observe that using prompts with richer information helps improve the repair effectiveness, which guides us to leverage static analysis tools in combination with Copilot Chat to enhance the security of automatically generated code. When using the fix command, security weaknesses in JavaScript code are resolved more frequently than those in Python, possibly because JavaScript code is easier to trace in context, allowing Copilot Chat to more readily identify and fix these issues. In contrast, Python often involves data structures, libraries, and system calls, which may make security weakness resolution less intuitive compared to JavaScript. However, when enhanced information is provided, the repair rates for security weaknesses in Python and JavaScript code segments do not differ significantly.
From the heatmap in Figure LABEL:fig:cwe-fix, we can see that the effectiveness of Copilot Chat in fixing different CWEs varies. Copilot Chat is particularly effective at fixing CWE-259: Use of Hard-coded Password and CWE-685: Function Call with Incorrect Number of Arguments, which pertain to access control issues and environment configuration. These issues are relatively straightforward and have standard repair methods, making them easier for automated tools to identify and fix. On the other hand, other weaknesses such as CWE-78: command injection and CWE-94: code injection involve executing commands or code provided by users, which typically require better complex contextual understanding and detailed input validation. Due to the complexity of these issues, Copilot Chat may encounter difficulties in automatically providing a fix. Additionally, we found that using more detailed instructions significantly improves repair outcomes for certain CWEs, such as CWE-330: Use of Insufficiently Random Values Weakness and CWE-79: Cross-site Scripting. This may be because our designed prompts provide richer information and context (the warning messages from static analysis tools), reducing ambiguity. This allows Copilot to effectively identify issues and generate correct fixes.
@@ 5.2. Implications
5.2.1. Importance of continuous security analysis of automatically-generated code
We conjecture that practitioners using Copilot and other AI code generation tools are likely to encounter security weaknesses in their generated code, regardless of the programming language used. Our study results reflect the inevitability of security weaknesses in automatically generated code. Practitioners must be aware of the diverse security scenarios in production and adopt multiple security prevention measures to address security risks before accepting vulnerable code suggestions.
Some CWEs, such as CWE-78: OS Command Injection and CWE-94: Code Injection, involve direct control of code execution by user inputs and may lead to severe security risks. Developers should be particularly vigilant about these high-risk operations and manually review these parts of the code. Input validation is the foundation of code security. Many security weaknesses (such as CWE-22: Path Traversal) arise from inadequate input validation. When using AI code generation tools, developers should avoid directly using user input as the basis for code execution or file path operations and manually add input validation or adopt a more secure coding approach. Moreover, Copilot performs relatively well in handling permission control (such as CWE-259: Use of Hard-coded Password) and environment configuration issues. Nevertheless, developers should be familiar with best practices related to permissions and verify that the configurations generated by AI code generation tools meet the project’s security requirements.
@@ 5.2.1. Importance of continuous security analysis of automatically-generated code
Maintaining a rigorous process of security analysis concurrently with code generation can help identify potential vulnerability issues and rectify the security weaknesses in time. Developers should follow the best practices for using code generation tools and always check the code suggestions generated by Copilot (or any code generation tools). For example, developers can establish a gated check-in build process to check and prevent security weaknesses when committing code generated by AI code generation tools. Initially, we can turn to automated tools to continuously scan the Copilot-generated code for known security weaknesses, such as CWEs. Then, Copilot Chat or other LLMs can be used to fix identified code with security weaknesses, followed by a code re-analysis with security analysis tools. At the same time, the severity of security weaknesses and the limitations of static analysis tools remind us that it is essential to conduct a subsequent manual assessment, including manual security code review for auto-generated code. By embracing this combined strategy for continuous security analysis, we can ensure a robust security shield for the code-committing process with AI-generated code.
5.2.2. Prevention of security weaknesses in AI-generated code
We offer the following suggestions on how to prevent potential security weaknesses in generated code:
Targeted security countermeasures: Based on the frequency of related CWEs in our dataset, practitioners can proactively prevent and address security weaknesses in a targeted manner. Developers should focus on dedicated CWEs in AI-generated code of different programming languages when using Copilot to generate e.g., Python or JavaScript code. Furthermore, a recent study shows that security weaknesses appear in certain code suggestions, but not all (Majdinasab et al., 2024); consequently, developers should carefully select potentially more secure suggestions, with the assistance of tools, that do not expose the code to vulnerabilities. Security weaknesses can vary across programming languages and application domains, and developers should pay special attention to these differences when using these automated code generation tools. For example, the security weaknesses path validation (such as CWE-22: Path Traversal) and dynamic execution (such as CWE-94: Code Injection) are more likely to occur in JavaScript code. At the same time, system calls and library loading in Python code are more likely to raise security concerns.
Standardized security assessment: Common security weaknesses in software development are also prevalent in the code generated by AI code generation tools. As a good practice, developers can use the CWE Top-25 list as a guide to understand which security weaknesses are most common and dangerous in the generated code. Additionally, the CWE Top-25 provides a standardized approach for security assessment, and developers can also use it as a guide to perform security audits of the AI-generated code. Using a standardized remediation strategy for common security weaknesses (such as privilege control and environment configuration) can effectively reduce the likelihood of generating insecure code. Developers can refer to the fixes provided in the static analysis tool documentation (CodeQL, 2024) or the mitigation measures for related CWEs provided by the MITRE (MITRE, 2024).
Enhancing prompt engineering to generate more secure code: The instruction tuning schemes in code generation not only impact the utility of the code but also its security. By combining the security fine-tuning with standard instruction tuning, joint optimization of security and utility can be promoted (He et al., 2024). It is recommended to incorporate security considerations from the initial stages of code generation. For specific security scenarios prone to CWEs, we can improve code security by enhancing prompt engineering, e.g., with the prompt patterns proposed in (White et al., 2023).
Distinguishing the complexity of security weaknesses: Some security issues, such as access control and environment configuration, have standard solutions and are well suited for quick fixes using tools like Copilot Chat. However, for complex logic controls (such as logic processing related to multi-layered inputs), Copilot Chat’s repair capabilities may be limited, requiring developers’ review and design. When using Copilot Chat, developers should assess the complexity of issues to determine which code can be automatically fixed and which requires manual intervention.
Static analysis tools combined with LLM-based chatbots for fixing security weaknesses: When using Copilot Chat to address security weaknesses, developers should provide more specific context and clearer instructions. More detailed information can significantly enhance Copilot Chat’s ability to fix security problems. Meanwhile, integrating static analysis tools allows for a more comprehensive identification of potential risks in the code. Developers can leverage static analysis tools to identify specific security risks and provide detailed instructions to Copilot Chat, enabling it to effectively fix the issues in Copilot-generated code.